OpenStudy (anonymous):

Can you find a number that lies one third of the distance between 1/3 and 2/3 ?

OpenStudy (anonymous):

@chycora

OpenStudy (anonymous):

I will give a medal

OpenStudy (dtan5457):

Trick question? lol

OpenStudy (anonymous):

I don't think so

OpenStudy (chycora):

ok i'm here

OpenStudy (dtan5457):

It's hard to get a number when the numerator and denominator is so small.. 1/3=3/9 2/3=6/9 difference of numerator =3 1/3 of 3=1 4/9?

OpenStudy (anonymous):

1/3=2/6 2/3=4/6 3/6 or 1/2

OpenStudy (chycora):

i think it's 2/9

OpenStudy (chycora):

wait 3/6 is the answer

OpenStudy (anonymous):

The real answer is 12/27 want to know how

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }*\frac{ 3 }{ 3 }=\frac{ 3 }{ 9 }\] \[\frac{ 2 }{ 3 }*\frac{ 3 }{ 3 }=\frac{ 6 }{ 9 }\] 

OpenStudy (chycora):

wat

OpenStudy (anonymous):

never mind continue

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }-\frac{ 1 }{ 3 }=\frac{ 2-1 }{ 3 }=\frac{ 1 }{ 3 }\]\[\frac{ 4 }{ 9 }-\frac{ 1 }{ 3 }=\frac{ 4-3 }{ 9 }=\frac{ 1 }{ 9 }=one~ third~of~\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

\[\frac{ 12 }{ 27 }~is~same~as~\frac{ 4 }{ 9 }\]

OpenStudy (anonymous):

basically you were right

OpenStudy (chycora):

ok yay yall gt the answer